Service Logbook¶
Date  | 
Action  | 
|---|---|
2025-03-21  | 
STELLA-2:  | 
2025-03-24  | 
STELLA-2: new holder for hydraulic oil-level switch  | 
2025-03-24  | 
STELLA-2: Fixed AZ-Limit guide (2-component glue)  | 
2025-03-21  | 
STELLA-2: Alt encoder cleaned  | 
2025-02-14  | 
STELLA-2: Electronics update (incomplete)  | 
2024-11-05  | 
STELLA-1: New hydraulic Oil filter  | 
2024-11-01  | 
WiFSIP: Overpressure valve at cold head sealed with cap & silicone  | 
2024-10-31  | 
WiFSIP: New hoses, new external filter, filled to 230 psi  | 
2024-10-31  | 
WiFSIP Cryotiger 6: added both pressure sensors  | 
2024-10-31  | 
WiFSIP CCD: unmounted & heated & pumped 24h  | 
2024-10-30  | 
sky: exchanged failed disk in slot 5  | 
2024-10-09  | 
ThAr1 HOK0933 out, HOK0934 in  | 
2024-10-09  | 
ThAr2 HLA1423 out, L74804 in  | 
2024-03-18  | 
WiFSIP Cryotiger 6 ca. 250 psi  | 
2024-03-18  | 
New firewall version & new fire computer online  | 
2024-03-18  | 
IP address of new guider1 set  | 
2024-03-15  | 
WiFSIPv2 in storage  | 
2024-03-15  | 
WiFSIPv3 pumped  | 
2024-03-15  | 
Communication boad in Controller changed  | 
2023-11-22  | 
New STELLA weather station (Thies)  | 
2023-03-03  | 
Fixed broken roof control modules  | 
2022-08-18  | 
ThAr1 HLA1421 out, HOK0933 in, Halo1 changed  | 
2022-08-18  | 
re-arranged energy chain WiFSIP  | 
2022-08-17  | 
WiFSIP Cryotiger 4 235 psi, new connector on supply side coldhead  | 
2022-08-16  | 
added some clearance (2/10mm) for calibration stage SES  | 
2022-08-15  | 
STELLA-1 / WiFSIP: Vacuum pumped  | 
2021-10-17  | 
ASIVA restarted  | 
2021-10-17  | 
removed supply pressure sensor from WiFSIP Cryotiger, added gas  | 
2021-10-16  | 
Update Becky-roof software  | 
2021-10-15  | 
installed Cryotiger 4 for Wifsip, added pressure sensors  | 
2021-10-15  | 
discarded Cryotiger 7 (leak)  | 
2021-10-14  | 
WiFSIP pumped  | 
2021-06-22  | 
STELLA-1 M3 cleaned  | 
2021-06-21  | 
STELLA-2 M1 cleaned  | 
2021-06-21  | 
STELLA-1 / WiFSIP: Cryotiger hoses (15m) and external filter exchanged  | 
2021-06-21  | 
STELLA-1 / WiFSIP: Vacuum pumped  | 
2021-06-21  | 
archive (161.72.132.71) finally replaced, old archive is now oldarchive  | 
2021-06-20  | 
filled Cryotiger 7 & replaced its internal filter, re-connected to WiFSIP  | 
2021-06-20  | 
filled Cryotiger 4 & replaced its internal filter  | 
2021-06-18  | 
archive: swapped SSD (drive 0)  | 
2021-05-18  | 
Firewall change (kept old SSD drive) (IAC, GREGOR)  | 
2021-03-23  | 
ThAr1 Nr HLA1421 in ( HHL1055 out) (IAC)  | 
2021-03-23  | 
ThAr2 Nr HLA1423 in ( HLA1422 out) (IAC)  | 
2020-12-14  | 
halo2 lamp change (IAC)  | 
2019-10-14  | 
SES shutter & calibration unit connected to serial3, serial2 removed  | 
2019-10-13  | 
STELLA-2 agunit connected to local network (gige2,guider2,serial2)  | 
2019-10-13  | 
new local network switch (Gbit)  | 
2019-10-13  | 
new sky computer, oldsky still online, setup of new archive/wifsip  | 
2019-10-12  | 
STELLA-1 old adapter disconnected (but aux1 still online)  | 
2019-10-11  | 
Gbit switch connected in lounge, cables cleaned up  | 
2019-10-11  | 
New STELLA-2 cable, but no change: module is broken  | 
2019-09-30  | 
WiFSIP CCD pumped Miguel/IAC  | 
2019-05-13  | 
WiFSIP CCD pumped  | 
2019-05-14  | 
Mirror cover drives STELLA1 lubricated  | 
2019-05-13  | 
Mirror cover hinches & drives STELLA2 repaired & lubricated  | 
2019-05-11  | 
Updated STELLA2-rackpc ntp startup script with ntpdate command  | 
2019-05-11  | 
created hwclock command & updated hwclock  | 
2019-05-10  | 
STELLA2-archive network cable exchanged  | 
2019-05-10  | 
new fiber network cable to STELLA2, connected as second uplink  | 
2019-05-10  | 
New bearings in most (exept most outward) roof guide wheels  | 
2019-05-09  | 
STELLA1 & STELLA2 hydraulic oil change, oil filter change  | 
2019-05-08  | 
Tightened roof wheels to specs (130Nm), drilled access holes for roof wheel blocks  | 
2019-02-26  | 
ASIVA lenses cleaned, taped temp-sensors cable.  | 
2019-02-26  | 
Halo2 lamp changed  | 
2019-02-25  | 
STELLA2 extraction pump exchanged phases (ran wrong way), cleaned up oil spill  | 
2019-02-15  | 
STELLA2 pump motor-cable changed to shielded (2 short pieces)  | 
2019-02-14  | 
STELLA2 pump motor-cable changed to shielded (long piece)  | 
2019-02-14  | 
STELLA1 M1/M3 cleaned  | 
2019-02-12  | 
STELLA2 M1 cleaned  | 
2019-02-12  | 
topped up Cryotiger 7, replaced WiFSIP external filter unit & female/female coupler  | 
2019-02-12  | 
connected Cryotiger #7 (S/N 1511CHE11116) to WIFSIP (instead of #6)  | 
2019-02-11  | 
STELLA1-focus-drive cleaned & lubricated  | 
2019-02-11  | 
STELLA1 short pump motor-cable changed to shielded  | 
2019-02-11  | 
connected Cryotiger #6 instead of 3, trashed 3  | 
2019-02-09  | 
re-connected dust sensor after repair  | 
2019-02-08  | 
WiFSIP new 15m flex hoses, filled gas, connected Cryotiger 3  | 
2018-11-27  | 
mittlere Längsführungsrollen kontrolliert & nachgezogen  | 
2018-11-27  | 
STELLA-1 pump motor-cable changed to shielded (but not the last 1m at the pump)  | 
2018-11-26  | 
vacuum WiFSIP  | 
2018-11-26  | 
fuse at ASIVA turned on again  | 
2018-11-26  | 
halo2 moved back a bit (same position as halo1)  | 
2018-11-26  | 
ThAr1 lamp changed new HHL1055 (old HHL0229)  | 
2018-10-30  | 
ses PE connections changed  | 
2018-10-30  | 
new version roof control software  | 
2018-10-29  | 
roof serviced (new limit switch positions  | 
2018-10-28  | 
softstarter statt motor relays in stella1 & 2 racks  | 
2018-10-25  | 
Dehumidifier connected  | 
2018-10-25  | 
WiFSIP CCD changed to new STA 4150A  | 
2018-10-25  | 
io3 & io1 moved to calibration switch (io5)  | 
2018-10-25  | 
STELLA roof control changed to individual limit switches  | 
2018-06-01  | 
Power supply for io1 & io2 replaced  | 
2018-04-13  | 
STELLA1 and STELLA2 level adjusted  | 
2018-04-13  | 
big A/C filter cleaned  | 
2018-04-12  | 
STELLA2 alt K80 changed to 1000 (was 2000 or 3000)  | 
2018-04-11  | 
roof drive gear oil changed  | 
2018-04-11  | 
ASIVA drive installed & started  | 
2018-01-22  | 
STELLA-1 replaced ETEL Az / Rob-X20 controller with Spare-Az  | 
2018-01-22  | 
STELLA-1 replaced ETEL Az controller with Rob-X20 & power supply  | 
2017-10-14  | 
removed ASIVA roof drive  | 
2017-07-11  | 
removed ASIVA roof drive  | 
2017-07-11  | 
refilled Cryotiger #4 (complete system refill, 225psi)  | 
2017-07-11  | 
WiFSIP new cryotiger hoses (just the two 15m hoses)  | 
2017-07-11  | 
WiFSIP vacuum  | 
2017-05-11  | 
SES vacuum  | 
2017-04-07  | 
WiFSIP vacuum  | 
2017-04-07  | 
ThAr2 Nr HLA1422 in (B68776 out)  | 
2017-04-06  | 
Baycam ethernet via Beckoff 100Mbit fiber switch, power via roof control  | 
2017-04-06  | 
Old-AGunit (STELLA-1, fiber feed) now power from PDU3 (also auxiliary1)  | 
2017-04-06  | 
New roof control, Beckhoff, weather station electronics  | 
2017-01-09  | 
switch SES calibration from ThAr2 to ThAr1  | 
2016-12-22  | 
ASIVA chain gear fixed  | 
2016-12-20  | 
SES & CCD with new CF-Cards.  | 
2016-12-20  | 
1Gbit uplink  | 
2016-12-20  | 
WiFSIP vacuum  | 
2016-12-19  | 
New firewall hardware, upgraded to leaf-Bering v6  | 
2016-12-17  | 
Debian 8 on sky, archive, wifsip  | 
2016-12-17  | 
Debian 8 on sky, archive, wifsip  | 
2016-08-02  | 
ASIVA cleaned & calibration run  | 
2016-08-01  | 
ASIVA chain gear fixed  | 
2016-08-01  | 
WiFSIP pressure gauge changed & pumped  | 
2016-07-31  | 
Weatherstation 2 - new brightness & additional rain2trigger sensors  | 
2016-07-30  | 
new WiFSIP network-fibre connection Gbit  | 
2016-07-09  | 
STELLA-1 M1 cleaned  | 
2016-07-08  | 
Halo1 & Halo2 changed (Daniel)  | 
2016-07-07  | 
STELLA-1 derotator controller exchanged  | 
2015-08-19  | 
changed ThAr1 (new Photon HHL0529, old HHL0530)  | 
2015-08-17  | 
pumped & baked WiFSIP dewar  | 
2015-08-17  | 
cleaned ASIVA lens, put seals to ASIVA rack, fixed gearbox, fixed sky computer  | 
2015-08-17  | 
cleaned M1 STELLA2  | 
2015-08-17  | 
modified roof control for manual control  | 
2015-08-16  | 
exchanged roof control modules 0C, 0E after lightning  | 
2015-08-16  | 
connected STELLA2 AG-unit via Gbit fiber directly to central ethernet switch  | 
2015-08-16  | 
replaced Cryotiger SES (Cryotiger #2) with new Cryotiger #5 (S/N L12432555)  | 
2015-08-16  | 
pressure before connecting 255psi, after connecting 235psi  | 
2015-08-15  | 
old pressure sensor back to SES, new seal, cleaned seal, pumped & baked SES dewar  | 
2015-08-14  | 
electrical power distribution cleaned up, new PDU4, new fuses  | 
2015-08-13  | 
connected Cryotiger #4 again (instead of #3)  | 
2015-08-13  | 
re-connected Filter unit, swapped hoses back to normal orientation, pressure 220psi  | 
2015-08-13  | 
purged the WiFSIP cold head with ISO-Propanol  | 
2015-07-28  | 
ThAr1 lamp defect, changed to ThAr2  | 
2015-07-08  | 
Rebooted STELLA1-tcc because landolt stars looked like donuts  | 
2015-06-24  | 
pumped SES  | 
2015-06-24  | 
pumped WiFSIP  | 
2015-06-24  | 
replaced SES-CCD pressure sensor & replaced signal cable  | 
2015-06-24  | 
removed filter unit on WiFSIP Cryosystem, switched supply/return on Cryotiger side  | 
2015-06-23  | 
replaced 1wire calib switches with moxa-io switch  | 
2015-06-23  | 
also on this moxa-io telescope 1 & 2 rack reset, lamp control  | 
2015-06-23  | 
attached Cryotiger #3, refilled Cryotiger #4, filter unit,15m hoses  | 
2015-06-23  | 
moved WiFSIP Cryotiger to E-room, added 5m hoses as extensions  | 
2015-06-23  | 
fixed network issue on STELLA1-aux guider cam, new replacement board in cupboard  | 
2015-05-06  | 
temporarily fixed broken ASIVA hatch mechanism  | 
2015-05-06  | 
replaced WiFSIP Cryotiger #4 (70psi) & hoses & filter with spares (Cryotiger #3, > 200psi)  | 
2015-04-08  | 
SES pumped remotely  | 
2015-02-23  | 
Pfeiffer vacuum pump attached to SES permanently, and remote controlled  | 
2015-02-21  | 
ThAr2 Nr B68776 in ( B00876 out)  | 
2015-02-19  | 
STELLA2 guider gige1 (instead of firewire cam)  | 
2014-07-02  | 
WiFSIP shutter/off-axis guider problems, reset of WiFSIP instrument and computer  | 
2014-06-20  | 
SES CCD vacuum pumped - IAC/Ignacio  | 
2014-06-18  | 
new UPS installed (Newave Upscale) - IAC  | 
2014-05-12  | 
(electronics room) A/C repaired  | 
2014-05-11  | 
STELLA-2 M1 cleaned reflectivity was 77/77/75/64, is 89/89/86/74  | 
2014-05-10  | 
STELLA-1 CCD centred a bit better, M2 aligned (less Coma)  | 
2014-05-10  | 
STELLA-1 M1, M2 and M3 cleaned  | 
2014-05-10  | 
STELLA-1 reflectivity was 75/69/71/69, 91/89/88/76, 83/82/81/68  | 
2014-05-10  | 
STELLA-1 reflectivity is (M1,2,3) 82/86/88/81, 91/91/89/77, 95/95/93/80  | 
2014-05-09  | 
VdN lightmeter cleaned  | 
2014-05-09  | 
ASIVA lenses cleaned  | 
2014-05-09  | 
ASIVA new chains in the hatch drive transmission  | 
2014-05-09  | 
Weatherstation 2 new T/H sensor, new analog Rain sensor  | 
2014-05-08  | 
SES readout window changed  | 
2014-05-06  | 
STELLA2 corrector cleaned  | 
2014-05-06  | 
SES pinhole change to 200mu  | 
2014-05-06  | 
SES dewar vacuum gepumpt  | 
2014-05-06  | 
SES slicer installed  | 
2013-12-17  | 
WiFSIP dewar vacuum gepumpt  | 
2013-12-17  | 
SES dewar vacuum gepumpt  | 
2013-12-17  | 
ThAr1 exchange S/N HHL0530 in, S/N B00863 out  | 
2013-12-16  | 
Dust monitor cleaned, filter changed, heater changed, fan changed, checked calibration.  | 
2013-12-02  | 
ThAr1 lamp defect, changed to ThAr2  | 
2013-08-15  | 
SES vacuum  | 
2013-08-15  | 
New CCD control computers (BOX PCs) with dedicated redundant DC power supplies  | 
2013-08-15  | 
WiFSIP dewar v2 installed, 2-channel CCD  | 
2013-07-19  | 
SES magellan controller new software version  | 
2017-07-16  | 
SES magellan controller new software version with communication problems -> useless images  | 
2013-04-17  | 
Using Halo-2 for flatfielding now  | 
2013-03-03  | 
SES dewar vacuum gepumpt  | 
2013-03-03  | 
Cryotiger #2 (SES) 210 psi  | 
2013-03-02  | 
SES shutter vom Tisch isoliert, shutterprogram geändert (ohne xma/xmd)  | 
2013-03-02  | 
STELLA1: M1 and M3 cleaned  | 
2013-03-02  | 
STELLA1: reflectivity was 53/49/51/48, 64/64/61/51  | 
2013-03-02  | 
STELLA1: reflectivity is (M1,2,3) 83/87/89/80, 91/91/89/76, 95/95/93/80  | 
2013-03-01  | 
calibration: Halo-lighhtbulb-positions moved away from fiber (to dim the light)  | 
2013-03-01  | 
STELLA2 tube balanced  | 
2013-03-01  | 
STELLA2-auxiliary mechanical alt-offset (1.5mm sheet metal on top feet)  | 
2013-03-01  | 
STELLA2: M1 cleaned: reflectivity (BGRIr) was 60/59/55/47, is 89/89/87/75  | 
2013-03-01  | 
SES: new fiber  | 
2013-02-08  | 
IPMI card connected to 192.168.1. network  | 
2013-02-07  | 
STELLA-1 parameters changed (K87) to decrease noise in SES CCD  | 
2013-02-06  | 
Temperature sensores lounge, lab installed and hooked into env database  | 
2013-02-05  | 
air condition filters cleaned (all 3 units)  | 
2013-02-04  | 
New Cryotiger #5 (S/N L12432555) in storage area  | 
2013-02-04  | 
wifsip disk 5 replaced  | 
2013-02-04  | 
wifsip & archive OS update, IMPI card installation  | 
2013-02-03  | 
sky OS update, IMPI card installation  | 
2013-02-03  | 
sky disk 15 replaced  | 
2013-02-03  | 
SES CCD controller: replaced fiber to the dsp board, now fiber exits through hole in box  | 
2013-01-07  | 
SES connection to CCD controller lost  | 
2012-12-20  | 
SES Heizung neu, mehr sensoren  | 
2012-12-19  | 
SES CCD format geändert, neuer server mit DATE-OBS  | 
2012-12-19  | 
SES faser ein& auskoppelung neu justiert  | 
2012-12-18  | 
SES dewar vacuum gepumpt  | 
2012-12-17  | 
WiFSIP dewar vacuum gepumpt  | 
2012-12-17  | 
Cryotiger #2 (SES): 210 psi  | 
2012-12-17  | 
Cryotiger #4 (WiFSIP): 200 psi  | 
2012-10-24  | 
11:25 Power meter 243752  | 
2012-10-22  | 
23:50 Power meter 243446  | 
2012-10-22  | 
Webcam2 (VTT) exchanged old Axis 200+ with Axis P1343  | 
2012-10-22  | 
SES exchange Thar2 lamp (new B00876, old B68775)  | 
2012-10-21  | 
heated & pumped SES dewar  | 
2012-10-21  | 
installed optocoupler for SES shutter, now new program on shutter controller  | 
2012-10-21  | 
disconnected rain sensor 2 (repair was not successful)  | 
2012-10-20  | 
STELLA-1 halved P-motor-current drive Parameters (no init possible)  | 
2012-10-19  | 
WiFSIP dewar pumped  | 
2012-10-19  | 
WiFSIP energy-chain split in two parts, cables held in the middle only  | 
2012-10-19  | 
Cryotiger #4 refilled (240 psi)  | 
2012-10-19  | 
WiFSIP Cryotiger #3 refilled (235psi), with new hoses,filter and head refilled  | 
2012-10-19  | 
WiFSIP Cryotiger pressure ca 220 psi  | 
2012-08-30  | 
SES shutter masse Leitungen verändert  | 
2012-08-29  | 
SES dewar neues Fenster/Feldlinse, vacuum  | 
2012-06-06  | 
WiFSIP dewar gepumpt  | 
2012-06-06  | 
SES neue optische & CCD camera  | 
2012-06-06  | 
SES neu justiert  | 
2012-03-19  | 
STELLA-2 rack Batterie gewechselt  | 
2012-03-19  | 
STELLA-1 neue Pollerverlängerungen  | 
2012-03-19  | 
STELLA-1 Strom-P-wert von 600 auf 400  | 
2012-03-17  | 
STELLA-1 derotator fixed holder limit-switch rocker  | 
2012-03-17  | 
vacuum WiFSIP CCD dewar  | 
2012-03-17  | 
WiFSIP computer changed disk 2 (p2)  | 
2012-03-15  | 
Cryotiger #4 190 psi  | 
2012-03-15  | 
WiFSIP repaired energy-chain mounting  | 
2012-03-15  | 
WiFSIP network cable repaired  | 
2012-03-14  | 
STELLA-1 derotator cable repaired (shortened, connector soldered)  | 
2011-12-10  | 
Wifsip neuer Halpha wide filter  | 
2011-12-08  | 
NO-motor hydraulic oil exchanged  | 
2011-12-04  | 
STELLA-1 hydraulic oil exchanged, hydraulic oil filter exchanged  | 
2011-12-03  | 
STELLA-2 hydraulic oil exchanged, hydraulic oil filter exchanged  | 
2011-12-01  | 
ASIVA re-aktiviert  | 
2011-11-30  | 
Windspeed-2 sensor exchanged  | 
2011-09-27  | 
Halo-1 working ?  | 
2011-09-27  | 
removed ASIVA optical camera and sent in for repair  | 
2011-09-27  | 
STELLA-1 M2-Spindel & Führungen gereinigt & geschmiert  | 
2011-09-26  | 
Linear stages AG-unit2 exchanged  | 
2011-09-26  | 
WiFSIP Pilar config changed to M3 mirror default position=2  | 
2011-09-25  | 
WiFSIP slighty changed energy-chain mounting on Az-Ring  | 
2011-09-25  | 
WiFSIP Cryotiger #4 (S/N LO8814101), with hoses, new filter,old (empty) coldhead  | 
2011-09-25  | 
WiFSIP pressure 202 psi  | 
2011-09-24  | 
WiFSIP dewar opened, closed, heated & pumped  | 
2011-09-24  | 
wifsip new computer  | 
2011-09-24  | 
ASIVA Lightmeter installed, new computer sky additionally connected via switch  | 
2011-09-23  | 
SES vacuum  | 
2011-08-25  | 
WiFSIP Cryotiger (nr 3?) offline (gas line pressure = 0dpi)  | 
2011-08-18  | 
using Halo-2 now, Halo-1 dead - wooz  | 
2011-07-13  | 
tele2 mediaconverter exchanged (died after a power cut on July 11 2011) - Ignacio  | 
2011-05-21  | 
new pinhole mirror (120 or 100µm)  | 
2011-05-21  | 
ThAr1 lamp exchanged (old S/N B00860, new B00863)  | 
2011-04-28  | 
ThAr1 lamp defect, changed to ThAr2  | 
2011-04-15  | 
Halo1 & Halo2 exchanged. Are now again 20W (but still on external power supplies).  | 
2011-04-15  | 
pinhole pos 335,308  | 
2011-04-15  | 
AG-2 new pinhole mirror (140µm), fiber adjusted, new calibration focus (65000)  | 
2011-04-15  | 
AG-2guider cam re-focused, new ADC-wheel offset, new exposure times for calibs  | 
2011-04-14  | 
WiFSIP CCD new pressure meter, fixed videoboard AD-3© problem  | 
2010-11-18  | 
AGunit-2 newly adjusted, fiber newly adjusted, new focuspositions etc.  | 
2010-11-16  | 
New PDUs for computers in Rack, new DC supplies in Rack  | 
2010-11-16  | 
STELLA-1 provisorische Streulichtblenden im Tubus  | 
2010-11-12  | 
WiFSIP CCD vacuum (for 3 days)  | 
2010-09-29  | 
Halo-1 lamp exchanged (IAC)  | 
2010-09-27  | 
Using Halo-2 for flatfielding now  | 
2010-09-13  | 
Halo-1 dark  | 
2010-08-21  | 
Auxiliary 1 & 2 cleaned  | 
2010-08-21  | 
ASIVA computer relocated to rack & power from UPS via 1wire switch  | 
2010-08-21  | 
STELLA-1 M3 cleaned  | 
2010-08-21  | 
STELLA-2 M1 cleaned  | 
2010-08-20  | 
STELLA-2 spider aligned  | 
2010-08-19  | 
SES neu justiert (fasereinkopplung, fokus, kollimatorausleuchtung)  | 
2010-08-19  | 
STELLA-2 Niveau korrigiert (war ca. 0.04mm off)  | 
2010-08-19  | 
STELLA-1 Niveau korrigiert (war ca. 0.06mm off)  | 
2010-08-19  | 
STELLA-1 PE-cable to AGunit-1 (old SES AGunit) changed(no connection to Az-ring any more)  | 
2010-08-18  | 
STELLA-1 oil extraction pump relay & motor exchanged  | 
2010-08-18  | 
WiFSIP power now through 1wire switch  | 
2010-08-17  | 
calibration light positions optimized  | 
2010-08-17  | 
Halo1 & Halo2 exchanged. Are now 50W.  | 
2010-05-11  | 
Fibers switched to STELLA-2 (old fibers still in place)  | 
2010-05-11  | 
ThAr2 lamp exchanged (old S/N B00857, new B68775)  | 
2010-04-14  | 
CryoTiger #3 (WiFSIP, S/N CM06241-0405) installed and connected, pressure is 215 psi  | 
2010-03-31  | 
ThAr2 lamp defect, changed to ThAr1  | 
2010-03-10  | 
replaced both Halogen lamps (without real need to…)  | 
2010-01-03  | 
Cryotiger #1 (SES, S/N CM05569-0903) replaced with Cryotiger #2 (SES, S/N CM05568-0903)  | 
2010-01-03  | 
Cryotiger SES Pressure is 225 psi  | 
2009-12-16  | 
SIGRIST dust monitor dismantled -> Spider invasion  | 
2009-12-16  | 
SES Korrektorplatte leicht gekippt -> Geister nun an hoffentlich günstigerer Position  | 
2009-12-15  | 
STELLA-2 Phasenwächter auf STELLA-1: 220 statt 230V, 30 statt 5 sec  | 
2009-12-15  | 
STELLA-1 motorkabel-Schirm von Stecker entkoppelt  | 
2009-12-14  | 
SES Faser neu justiert  | 
2009-12-14  | 
STELLA-1 M1,M3 cleaned  | 
2009-12-14  | 
AG-unit-1 Seitenwechsel - Fokus sollte nun -10mm sein  | 
2009-12-14  | 
Derotator Teleskop I, Seitenwechsel  | 
2009-06-30  | 
E-Raum Klimaanlage Filter gereinigt  | 
2009-06-30  | 
Drehstifte der großen Klappen v. STELLA-I abgeschliffen & geschmiert  | 
2009-06-30  | 
Softwareupdate STELLA-II (M. Ruder)  | 
2009-06-26  | 
Softwareupdate STELLA-I (M. Ruder)  | 
2009-06-26  | 
Kabel X12 und X6, STELLA-II getauscht  | 
2009-06-25  | 
Kabel X12 und X6, STELLA-I getauscht  | 
2009-05-28  | 
STELLA-1 Fokus-Führungen gereinigt  | 
2009-05-28  | 
STELLA-1 defekte Ader Ölrücklaufpumpe überbrückt  | 
2009-05-28  | 
STELLA-1 Scharnier Spiegelklappe klein geschmiert/geschliffen  | 
2009-05-28  | 
STELLA-1 Öl nachgefüllt (nach Verlust großer Mengen)  | 
2009-04-24  | 
STELLA-2 corrector position corrected with 3.5mm Spacers  | 
2009-04-24  | 
STELLA-1 M2 drive cleaned  | 
2009-04-24  | 
temp[0] = -118.001373 2780  | 
2009-04-23  | 
STELLA-1 M1,M3 cleaned  | 
2009-04-21  | 
SIGRIST dust monitor installed  | 
2009-03-02  | 
STELLA-2 Niveau korrigiert (war ca. 0.08mm off)  | 
2009-03-01  | 
Entfeuchtungs-Granulat SES ausgeheizt  | 
2009-03-01  | 
SES cover geöffnet (AKF gesucht)  | 
2009-03-01  | 
Klimaanlage E-Raum Filter gereinigt  | 
2009-03-01  | 
STELLA-1 Niveau korrigiert (war einige 0.1mm off)  | 
2009-02-28  | 
STELLA-1 Azimut ETEL controller exchange (defect)  | 
2008-11-02  | 
Cleaned A/C filter & Dehumidifier filter  | 
2008-11-02  | 
ThAr1 exchange S/N B0080 in, S/N B00220 out  | 
2008-10-31  | 
Entfeuchtungs-Granulat SES ausgeheizt  | 
2008-10-29  | 
SES dewar temperature back to normal (-188/-120)  | 
2008-10-27  | 
SES vacuum pumped, Cryotiger pressure is 210 psi  | 
2008-10-09  | 
ThAr1 lamp defect, changed to ThAr2  | 
2008-07-31  | 
STELLA-2 Niveau korrigiert (war ca 0.1mm off)  | 
2008-07-30  | 
STELLA-2 Hydraulik-Oel nachgefuellt (ca 40 Liter)  | 
2008-07-30  | 
STELLA-2 Kabel getauscht (Kabel wieder defekt)  | 
2008-06-02  | 
STELLA-2 Kabelbruch repariert (ungenutzte Ader statt gebrochener verwendet)  | 
2008-06-02  | 
Klimaanlage E-Raum Filter gereinigt  | 
2007-11-21  | 
Optimized calib-light alignment in CalibUnit and AGunit  | 
2007-11-21  | 
New exposure times are 13 sec for both flat lamps and 20 sec for both ThAr  | 
2007-11-21  | 
Wasser in UPS nachgefuellt  | 
2007-11-21  | 
CalibUnit neu justiert, Motor Parameter geaendert  | 
2007-11-20  | 
ThAr Netzteile nun mit Strom-Ueberwachung versehen  | 
2007-11-20  | 
ThAr2 Lampe getauscht  | 
2007-11-20  | 
ThAr-Lampennetzteil 2 getauscht (das alte nr.1)  | 
2007-08-24  | 
Klimaanlage E-Raum Filter-counter reset  | 
2007-08-23  | 
Cryotiger not at correct temperature  | 
2007-08-23  | 
Ethernet wire from archive to TelescopeRack1 (to run Pilar on archive)  | 
2007-08-23  | 
Update archive to Debian 4.0/ Java 1.5,downgrade to 1.4 because of xml problems  | 
2007-08-22  | 
Power to Cryo etc. (see reset switch) now from Gallaxy UPS  | 
2007-08-22  | 
Reset switch for Cryotiger/Allsky/shutter/calibration unit installed  | 
2007-08-22  | 
Reset switch of AGunit correctly wired  | 
2007-08-22  | 
ETEL Controller Alt STELLA 2 Motor-current-P parameter decreased  | 
2007-08-21  | 
ETEL Controller Alt STELLA 2 exchanged  | 
2007-08-20  | 
ThAr Power supply connections checked  | 
2007-07-24  | 
ETEL Controller Alt STELLA 2 defekt, zur Reparatur nach D  | 
2007-07-24  | 
Cryotiger at normal Temp. again  | 
2007-07-20  | 
CUO CCD pumped  | 
2007-07-19  | 
ThAr-Lampennetzteil 1 getauscht  | 
2007-07-19  | 
ThAr1 & ThAr2 Lampen getauscht  | 
2007-07-17  | 
Klimaanlage E-Raum Filter gereinigt  | 
2007-03-30  | 
ThAr2 defekt, stellen auf ThAr1 um  | 
2007-03-21  | 
STELLA-I port 2&3 vertauscht (kabel lt. Schaltplan-Aenderung getauscht)  | 
2007-03-21  | 
SES cryotiger pressure 210 psi  | 
2007-03-21  | 
SES cryotiger cleaned  | 
2007-03-21  | 
CUO CCD pumped  | 
2007-03-20  | 
STELLA-1 software update  | 
2007-03-20  | 
STELLA-2 encoder tuned  | 
2006-12-14  | 
AG-I Beam splitter flipped  | 
2006-12-14  | 
AG-I & SES Fiber coupling adjusted  | 
2006-12-14  | 
Fan exchange & air filter exchange on master/ses/fire  | 
2006-09-09  | 
STELLA-I M1 and M-3 cleaned  | 
2006-01-31  | 
SES cryotiger pressure 210 psi  |